please be as clear as possible, take note of units and significant figures. thanks for the helpAn engineer performed N= 20 tests to assess the load capacity of a new anchoring device. The measured load capacities are as
Answer #1

N = 20

Let x be the load capacities

x

(x-\bar{x})^2

=(x-19.1)^2

11.0 65.61
20.0 0.81
24.0 24.01
13.0 37.21
16.0 9.61
24.0 24.01
18.0 1.21
7.0 146.41
20.0 0.81
22.0 8.41
17.0 4.41
18.0 1.21
28.0 79.21
27.0 62.41
20.0 0.81
15.0 16.81
12.0 50.41
11.0 65.61
31.0 141.61
28.0 79.21
\sum (x-\bar{x})^2=819.8

Q.1.

a)

Mean of the population:

\mu=\frac{\sum x}{N}=\frac{382.0}{20}=19.1

Mean of the population is 19.1

Standard deviation of the population:

σ Σ(1 – 1)2 N

\sigma=\sqrt{\frac{819.8}{20}}

\sigma=\sqrt{40.99}

\sigma=6.402 (Round to 3 decimal)

Standard deviation of the population is 6.402

Coefficient of variation (CV) :

CV=\frac{\sigma}{\mu}*100

CV=\frac{6.402}{19.1}*100

CV=0.3352*100

CV = 33.52%

Coefficient of variation is 33.52%

Q.2

b)
n = sample size = 9

Consider first 9 values from above 20 values as a sample.

x

(x-\bar{x})^2

=(x-17)^2

11.0 36
20.0 9
24.0 49
13.0 16
16.0 1
24.0 49
18.0 1
7.0 100
20.0 9
\sum x=153.0 \sum (x-\bar{x})^2=270

Point estimate of mean:

\bar{x}=\frac{\sum x}{n}=\frac{153.0}{9}=17

Point estimate of mean is 17

Point estimate of Standard deviation:

s=\sqrt{\frac{\sum (x-\bar{x})^2}{n-1}}

s=\sqrt{\frac{270}{9-1}}

s=\sqrt{\frac{270}{8}}

s=\sqrt{33.75}

s=5.809 (Round to 3 decimal)

Point estimate of Standard deviation is 5.809

Point estimate of Coefficient of variation (CV) :

CV=\frac{s}{\bar{x}}*100

CV=\frac{5.809}{17}*100

CV=0.3417*100

CV = 34.17%

Point estimate of Coefficient of variation is 34.17%

c)
Standard error of the mean:

SE=\frac{s}{\sqrt{n}}

SE=\frac{5.809}{\sqrt{9}}

SE=\frac{5.809}{3}

SE = 1.936    (Round to 3 decimal)

Standard error of the mean is 1.936

d)
90% confidence interval :

COnfidence level = c = 0.90

Sample size = n = 9

Sample mean = T = 17

Population standard deviation = \sigma = 6.402

Here Population standard deviation is known so we use z interval.

90% confidence interval is

\bar{x}-z_c*\frac{\sigma}{\sqrt{n}}<\mu<\bar{x}+z_c*\frac{\sigma}{\sqrt{n}}

where zc is z critical value for (1+c)/2 = (1+0.90)/2 = 0.95

zc = 1.645 (From statistical table of z values, average of 1.64 and 1.65,(1.64+1.65)/2 = 1.645)

17-1.645*\frac{6.402}{\sqrt{9}}<\mu<17+1.645*\frac{6.402}{\sqrt{9}}

17-1.645*2.134<\mu<17+1.645*2.134

17-3.51043<\mu<17+3.51043

13.490<\mu<20.510 (Round to 3 decimal)

90% confidence interval is (13.490, 20.510)